(25 PTS) 3. Solve the given system that satisfied the given initial condition. X(t) = 12° - 1]xce), x(0) = 1;}]: - (25 p
Answer #1

3. X(t) = Act) X(0) -3 -1 2~1 0 Now, A= - 3 - 2 -1 det (A-AI) -3-1- 2. 1-X -(3-X)(-1-) - (2)(-1) x2 + 4x +3+2 x2 + 40+ 5 (2+Page for ^2=-2t & A-AI --1 -1- 2. 1 1-1 2. 2 1 1 O V. JE 9 V4 = t (1+i + (1+i) , V = t 2 ie, t= - ₂ + 2 1 = X(t)=0; ekz-E-419= x(+) = Cz 0-2t -cost-sint 2 cost .t4 e 2t -cost tsint - sint X(o) = (-43-4) 11 2 => C3 = 0 W (4=2 X(t) = 2 0-24 (-cost tsin